The following information concerns the intangible assets of Epstein Corporation:

  1. On June 30, 2021, Epstein completed the acquisition of the Johnstone Corporation for $1,460,000 in cash. The fair value of the net identifiable assets of Johnstone was $1,250,000.
  2. Included in the assets purchased from Johnstone was a patent that was valued at $73,800. The remaining legal life of the patent was 14 years, but Epstein believes that the patent will only be useful for another nine years.
  3. Epstein acquired a franchise on October 1, 2021, by paying an initial franchise fee of $180,400. The contractual life of the franchise is 11 years.


Required:

1. Prepare year-end adjusting journal entries to record amortization expense on the intangibles at December 31, 2021.
2. Prepare the intangible asset section of the December 31, 2021, balance sheet

Partial Balance Sheet
December 31, 2021
Intangible assets:
Goodwill
Patent
Franchise
Total intangibles
Answer #1
1
Debit Credit
December 31, 2021 NO journal entry required
December 31, 2021 Amortization expense 4100 =73800/9*6/12
      Patent 4100
December 31, 2021 Amortization expense 4100 =180400/11*3/12
    Franchise 4100
2
Partial Balance Sheet
December 31, 2021
Intangible assets:
Goodwill 210000 =1460000-1250000
Patent 69700 =73800-4100
Franchise 176300 =180400-4100
Total intangibles 456000